# Wireless Datasets for Aerial Networks

### v1

@@ VIII-D Possible Uses of Dataset
Use cases of this data set include:
VIII-D1 LoRaWAN Optimization
Analyze gateway performance to optimize deployment strategies for Internet of Things (IoT) networks.
VIII-D2 Data Reliability Studies
Assess the reliability of transmitted data under varying conditions (e.g., noise levels, signal strength).
VIII-D3 IoT-Based Navigation Systems
Support development of IoT-enabled vehicle navigation and telemetry systems.
VIII-D4 Energy Efficiency Analysis
Explore power consumption patterns in LoRaWAN devices.
@@ IX-A Description of Hardware and Software
((a))
((b))
((c))
Figure 21: Hardware and system architecture used for A2G multipath channel sounding measurements. (a) Field deployment of the portable measurement node based on USRP B210, mounted on a UAV platform for aerial data collection. (b) Custom-designed GNSSDO providing high-precision timing and synchronization. (c) Block diagram of the overall channel sounder system, highlighting key components and signal flow used during the A2G experiments [44].
The experimental setup involved a UAV and a fixed node located at the AERPAW Lake Wheeler testbed site in Raleigh, North Carolina. The UAV carries a portable node equipped with a USRP B210, RF front-end, Intel NUC, and a custom-designed Global Navigation Satellite System disciplined oscillator (GNSSDO), as shown in Fig. 21(a). The fixed node is configured with identical equipment.
Accurate characterization of A2G multipath wireless channels requires precise temporal alignment between the transmitting and receiving devices. Multipath propagation inherently introduces time-varying delays and phase shifts, and without stringent synchronization, these effects can be misinterpreted as genuine channel behavior. The custom-designed GNSSDO system, as shown in Fig. 21(b), employed by AERPAW ensures a stable and common timebase for both the UAV and fixed node, minimizing phase ambiguity and enabling reliable quantification of multipath components and time-domain channel impulse responses (CIRs). The GNSSDO system achieves 2.5 ns pulse-per-second (PPS) accuracy between the nodes, facilitating high-fidelity channel measurements.
Fig. 21(c) presents the overall architecture of the channel sounder software. The software, developed in Cython, enables real-time processing of the USRP B210’s full bandwidth (56 MHz) and beyond, as described in [43]. The USRP B210 is synchronized using GNSSDO-generated PPS and 10 MHz reference signals, ensuring accurate timing control and reliable data acquisition. The system supports multiple configurable sounding waveforms, including Zadoff-Chu, pseudo-noise (PN), and chirp sequences. For the measurements in this study, a Zadoff-Chu sequence of length 401 and root index 200 was used, with each sequence repeated four times. The sounding was conducted at a measurement frequency of 4 Hz, with a center frequency of 3564, 3620, or 3686 MHz, and a transmit power of 19 dBm. A total of nine flight experiments were performed at three altitudes (30, 60, and 90 meters) along a 500-meter flight path, with a flight speed of 5 m/s. This configuration supports a wide range of channel sounding scenarios with high temporal and spatial resolution. For additional information regarding software, please refer to [43].
@@ IX-B Dataset Format
The raw channel sounding data is stored in SigMF [64]. Each dataset consists of two files:
  • .sigmf-data: a binary file containing complex I/Q samples for a given timestamp in 32-bit float format (little-endian).
  • .sigmf-meta: a JSON file that provides metadata, including sample rate, center frequency, timestamp, and capture details such as the transmit waveform and sequence configuration.
The metadata file follows the SigMF specification and includes additional custom fields to describe:
  • GPS coordinates and altitude of the UAV during each capture,
  • sequence length and root index of the Zadoff-Chu waveform,
  • measurement frequency and timestamp synchronization parameters,
  • flight and experimental configuration details.
((a))
((b))
Figure 22: Representative results from the A2G channel sounding campaign using a UAV-mounted transmitter. (a) Representative CIR illustrating multipath characteristics observed during flight. (b) Corresponding path loss measurement as a function of time, recorded at an altitude of 90 meters [44].
This format ensures reproducibility, interoperability with third-party tools, and ease of integration with standard signal processing pipelines. The GitHub repository also provides post-processing software, including Python scripts and a Jupyter notebook that operate on either the compressed .npz format or directly on the .sigmf-data files. Users can reproduce the published figures and analysis by running the provided PostProcess.ipynb notebook, which implements correlation-based channel response extraction and other visualization tasks.
@@ IX-C Representative Results
This section presents representative results from channel sounding experiments conducted at the AERPAW LWRFL, a rural environment designed for wireless experimentation. Fig. 22(a) illustrates a representative CIR obtained using correlation-based processing of the received Zadoff-Chu sequences. The three-dimensional plot shows the evolution of multipath components over time during a UAV flight. The horizontal axis represents the delay (in microseconds), the vertical axis indicates the experiment time (indexed per snapshot), and the color scale denotes the received signal magnitude in dB. The black trace projected onto the back plane indicates the UAV’s GPS distance from the fixed transmitter as a function of time. The variation in delay spread and path power over time reflects the dynamic nature of the A2G propagation channel, including the impact of UAV motion and altitude variation. Fig. 22(b) shows the corresponding path loss measurement over the full flight trajectory. The UAV’s altitude profile is overlaid to highlight different flight phases, including takeoff, flight, and landing. The results demonstrate a clear relationship between the received power and UAV position, consistent with expected large-scale path loss behavior.
@@ IX-D Possible Uses of Dataset
IX-D1 A2G Channel Modeling
Researchers can extract key propagation characteristics such as delay spread, Doppler profiles, path loss exponents, and coherence bandwidth to develop statistical or geometry-based A2G channel models.
IX-D2 Machine Learning for Wireless
The dataset can be used to train and evaluate machine learning models for tasks such as channel estimation or link quality prediction in UAV-based networks.
IX-D3 Impact of Altitude and Mobility
With measurements conducted at multiple altitudes and along a defined flight trajectory, the dataset supports analysis of how UAV height and speed affect propagation conditions and coverage.
IX-D4 Waveform and system Design
The raw I/Q data and metadata can be used to simulate and evaluate the performance of waveform designs under realistic conditions.
IX-D5 Validation of Ray-Tracing or Analytical Models
The dataset provides empirical benchmarks that can be used to validate and calibrate ray-tracing simulations or analytical propagation models in rural environments.
@@ X Wireless Localization Dataset
((a))
((b))
((c))
Figure 23: Overview of the AERPAW testbed showing the RF sensor, UAV node, and LWRFL site layout [48].
@@ X-A Description of Hardware and Software
UAV flight experiments were conducted in AERPAW’s controlled environment to generate both datasets, using Keysight N6841A RF sensors to collect TDOA measurements. The N6841A (Fig. 23(a)) is capable of detecting, recording, and precisely time-stamping RF signals across a frequency range of 20 MHz to 6 GHz, with a maximum bandwidth of 20 MHz. Equipped with a broadband omnidirectional antenna and GPS-based timestamping, the system uses Keysight’s N6854A Geolocation Software and Sensor Management Tool to support TDOA, RSS, and hybrid localization methods for tracking RF sources within approximately a 2 km radius. However, the software is limited to 2D localization and does not estimate altitude.
Fig. 23(c) shows the deployment of these sensors at the LWRFL, where a single N6841A unit is mounted on each of the four towers labeled LW2 through LW5, approximately 10 meters above ground level. The sensors are deployed in a rural environment with mixed LoS conditions due to tree cover and building obstructions. All RF sensors within AERPAW are synchronized using a shared GPS-disciplined clock infrastructure and operate on the same local network, with individual IP addresses assigned within a common subnet. Centralized management via the Keysight Geolocation Server ensures network-level time synchronization, which is essential for accurate TDOA-based localization. During experiments, the N6841A units capture I/Q data from RF signals, which are subsequently processed to estimate UAV positions. However, the specific algorithms used for TDOA extraction and position estimation are proprietary to Keysight and not publicly available.
In our previous work [48], which focused on evaluating TDOA-based UAV localization accuracy under varying altitude and signal bandwidth conditions, measurements were compared to the Cramér-Rao Lower Bound (CRLB) to assess localization performance limits. In this study, a 3.32 GHz channel sounding waveform was transmitted from a UAV-mounted SDR (Fig. 23(b)) and RF sensors localized the UAV along multiple repeated trajectories. The dataset associated with this study is publicly available at [45]. Similarly, in a separate study by our group [47], which proposed an extended Kalman filter (EKF) framework to improve UAV tracking performance, TDOA-based localization was performed using downlink control signals from a DJI Inspire 2 UAV operating in the 2.400-2.483 GHz ISM band with a 20 MHz bandwidth. The dataset corresponding to this experiment is also publicly available [46].
@@ X-B Dataset Format
The dataset associated with [48, 45] is provided in the file TDOA_UAV_Localization_Data.zip. This archive contains one CSV file for each flight, with the altitude, signal bandwidth, and recording date indicated in each filename. For example, the file 40m_1.25MHz_7.15.24.csv corresponds to a flight conducted at an altitude of 40 meters, using a 1.25 MHz signal bandwidth, on July 15th, 2024. The archive includes a total of five CSV files, each representing an individual flight. Each file contains the following recorded information:
  • Center frequency of the signal.
  • Latitude and longitude estimates for the UAV.
  • Ground truth coordinates (latitude, longitude, and altitude) labeled with “GT”.
  • RHO (degree of cross-correlation between received signals ranging from 0 to 1, where 1 indicates perfect correlation) and CEP (circular error probability) metrics for localization performance.
  • Timestamps corresponding to each position estimate.
  • A binary variable LOStoLW2-5 indicating LoS status to AERPAW towers LW2 through LW5:
    • A value of 1 indicates LoS to the respective tower.
    • A value of 0 indicates NLoS to the respective tower.
If both the latitude and longitude values for a given measurement are recorded as 0, this indicates that the geolocation software failed to produce a valid position estimate. The dataset archive also includes a MATLAB helper script, KeysightRTDOALocalizationforFlights.m, which assists in processing TDOA-based UAV localization data from a selected CSV file and generates visualizations and performance metrics for evaluating UAV position estimation accuracy.
The dataset associated with [47, 46] follows the same general format but differs in two key ways. First, the dataset for [47, 46] does not include LoS indicator variables. Second, the data is split into two separate CSV files: Inspiron_backup.csv, containing the field measurements collected by the Keysight RF sensor, and GPS_Flight1_backup.csv, containing the ground truth trajectory data recorded using a GPS application.
@@ X-C Representative Results
((a))
((b))
@@
Fig. 24(a) illustrates the UAV’s 40-meter altitude trajectory overlaid on the LWRFL site map, showing LoS conditions relative to three RF sensor towers (LW2, LW3, and LW4), which are indicated by green triangles. The trajectory is color-coded based on five distinct LoS scenarios, where each case indicates LoS or NLoS status relative to the three towers. For example, Case 1 (green) represents segments where the UAV maintained LoS to all three towers, while Case 5 (purple) corresponds to areas where only LW3 maintained LoS. The color-coding of the trajectory clearly highlights transitions between LoS conditions as the UAV moves through the field, illustrating the mixed LoS/NLoS environment characteristic of the LWRFL site. This visualization provides important context for interpreting localization performance, as LoS availability directly influences TDOA measurement quality and positioning accuracy.
Fig. 24(b) presents a color-coded visualization of localization error along the UAV’s trajectory. Each point represents a ground truth position of the UAV, with circles indicating the corresponding localization error relative to the nearest estimated coordinate. Green circles mark locations where the distance error is less than 100 meters, while red circles identify points where the error exceeds 100 meters. Blue triangles indicate the locations of the RF sensor towers used for TDOA-based localization.
As shown in the figure, the majority of the trajectory is associated with low localization error (green), suggesting consistent and accurate TDOA-based position estimates along most of the UAV’s flight path. Higher error regions (red) are concentrated near certain trajectory segments corresponding to NLoS conditions or poor geometric dilution of precision relative to the sensor towers. This visualization highlights how mixed LoS/NLoS conditions and sensor geometry influence localization performance throughout the trajectory.
@@ X-D Possible Uses of Dataset
X-D1 Tracking Filter Development/Evaluation
The dataset can support the development and evaluation of tracking filters, such as Kalman or particle filters, by providing real-world UAV trajectory and localization measurement sequences for state estimation under mixed LoS/NLoS conditions.
X-D2 TDOA Localization Algorithm Benchmarking
Supports performance evaluation and comparison against Keysight’s industry-standard TDOA-based localization methods using real-world RF measurements collected under mixed LoS/NLoS conditions.
X-D3 Sensor Fusion
Enables the development and evaluation of multi-sensor fusion algorithms by combining TDOA-based RF localization data with complementary modalities such as inertial, visual, or radar measurements for improved UAV tracking and state estimation.
X-D4 NLoS Modeling and Analysis
Supports characterization and modeling of NLoS conditions by providing real-world localization errors and LoS/NLoS labels, enabling research into NLoS detection, mitigation, and bias-aware localization techniques.
X-D5 CRLB Modeling and Validation
Enables theoretical performance analysis through CRLB modeling and supports empirical validation by comparing theoretical bounds against real-world TDOA-based localization errors under varying sensor geometries and environmental conditions.

### v2

@@ VIII-D Possible Uses of Dataset
The use cases of this data set include:
LoRaWAN Optimization for Mobile Aerial Applications: The dataset enables evaluation of LoRaWAN performance in both urban and rural environments under mobile and aerial operating conditions. By analyzing received signal strength, packet delivery behavior, and gateway visibility as functions of UAV position and altitude, researchers can assess coverage limitations and inform gateway placement and density strategies. These measurements are particularly relevant for IoT deployments involving mobile or airborne platforms, which remain comparatively underrepresented in existing LoRaWAN measurement studies.
LoRa Physical-Layer Characterization: The dataset supports empirical analysis of LoRa physical-layer behavior under varying channel conditions, including changes in noise floor and received signal strength. The availability of altitude- and orientation-tagged measurements enables characterization of air-to-ground LoRa propagation trends and transmission reliability. Such data can be used to evaluate physical-layer robustness and to inform propagation modeling efforts for low-power wide-area aerial communication scenarios.
@@ IX-A Description of Hardware and Software
The experimental setup involved a UAV and a fixed node located at the AERPAW Lake Wheeler testbed site in Raleigh, North Carolina. The UAV carries a portable node equipped with a USRP B210, RF front-end, Intel NUC, and a custom-designed Global Navigation Satellite System disciplined oscillator (GNSSDO), as shown in Fig. 20a. The fixed node is configured with identical equipment.
Accurate characterization of A2G multipath wireless channels requires precise temporal alignment between the transmitting and receiving devices. Multipath propagation inherently introduces time-varying delays and phase shifts, and without stringent synchronization, these effects can be misinterpreted as genuine channel behavior. The custom-designed GNSSDO system, as shown in Fig. 20b, employed by AERPAW ensures a stable and common timebase for both the UAV and fixed node, minimizing phase ambiguity and enabling reliable quantification of multipath components and time-domain channel impulse responses (CIRs). The GNSSDO system achieves 2.5 ns pulse-per-second (PPS) accuracy between the nodes, facilitating high-fidelity channel measurements.
(a) Experiment setup for the USRP B210-equipped portable node and UAV.
(b) Custom-designed GNSSDO system.
(c) Overall system architecture of the channel sounder.
Figure 20: Hardware and system architecture used for A2G multipath channel sounding measurements. (a) Field deployment of the portable measurement node based on a USRP B210, mounted on a UAV platform for aerial data collection. (b) Custom-designed GNSSDO providing high-precision timing and synchronization. (c) Block diagram of the channel sounder, highlighting key components and signal flow used during the A2G experiments.
Fig. 20c presents the overall architecture of the channel sounder software. The software, developed in Cython, enables real-time processing of the USRP B210’s full bandwidth (56 MHz) and beyond, as described in [35]. The USRP B210 is synchronized using GNSSDO-generated PPS and 10 MHz reference signals, ensuring accurate timing control and reliable data acquisition. The system supports multiple configurable sounding waveforms, including Zadoff-Chu, pseudo-noise (PN), and chirp sequences. For the measurements in this study, a Zadoff-Chu sequence of length 401 and root index 200 was used, with each sequence repeated four times. The sounding was conducted at a measurement frequency of 4 Hz, with a center frequency of 3564, 3620, or 3686 MHz, and a transmit power of 19 dBm. A total of nine flight experiments were performed at three altitudes (30, 60, and 90 meters) along a 500-meter flight path, with a flight speed of 5 m/s. This configuration supports a wide range of channel sounding scenarios with high temporal and spatial resolution.
@@ IX-B Dataset Format
The A2G channel sounding dataset is released using the SigMF specification [85], enabling standardized representation of raw signal measurements together with comprehensive metadata. The dataset captures synchronized complex baseband samples acquired during UAV-based channel sounding experiments and supports detailed analysis of multipath propagation, path loss, and Doppler effects.
The associated metadata describe measurement parameters such as sampling rate, center frequency, capture timing, and waveform configuration, along with UAV state information including geographic location, altitude, and experimental context. Additional fields document the Zadoff–Chu sequence parameters and synchronization settings used during the sounding process, enabling reproducible extraction of channel impulse responses and derived propagation metrics.
Representative results obtained from the dataset are illustrated in Fig. 21, including an example channel impulse response and corresponding path loss measurements collected during a UAV flight at 90 m altitude. These examples demonstrate the suitability of the dataset for multipath characterization and propagation modeling.
(a) Representative CIR from channel sounding experiment.
(b) Representative path loss measurement from channel sounding experiment.
Figure 21: Representative results from the A2G channel sounding campaign using a UAV-mounted transmitter. (a) Representative CIR illustrating channel impulse response characteristics observed during flight. (b) Corresponding path loss measurement as a function of time, recorded at an altitude of 90 meters.
The dataset is accompanied by post-processing tools that support channel response extraction and visualization using standard signal processing workflows. Detailed file structure, binary format descriptions, and post-processing utilities are documented in Appendix A. This organization ensures interoperability with third-party tools while preserving full reproducibility.
@@ IX-C Representative Results
This section presents representative results from channel sounding experiments conducted at the AERPAW LWRFL, a rural environment designed for wireless experimentation. Fig. 21a illustrates a representative CIR obtained using correlation-based processing of the received Zadoff-Chu sequences. The three-dimensional plot shows the evolution of multipath components over time during a UAV flight. The horizontal axis represents the delay (in microseconds), the vertical axis indicates the experiment time (indexed per snapshot), and the color scale denotes the received signal magnitude in dB. The black trace projected onto the back plane indicates the UAV’s GPS distance from the fixed transmitter as a function of time. The variation in delay spread and path power over time reflects the dynamic nature of the A2G propagation channel, including the impact of UAV motion and altitude variation. Fig. 21b shows the corresponding path loss measurement over the full flight trajectory. The UAV’s altitude profile is overlaid to highlight different flight phases, including takeoff, flight, and landing. The results demonstrate a clear relationship between the received power and UAV position, consistent with expected large-scale path loss behavior.
@@ IX-D Possible Uses of Dataset
Air-to-Ground Channel Modeling: The dataset enables extraction of key air-to-ground propagation characteristics, including delay spread, Doppler profiles, path loss behavior, and coherence bandwidth. These measurements support the development and evaluation of statistical and geometry-based A2G channel models grounded in empirical observations.
Machine Learning for Wireless Systems: The dataset supports training and evaluation of machine learning models for UAV-based wireless applications. The availability of synchronized signal measurements and metadata enables data-driven approaches to tasks such as channel estimation and link quality prediction under realistic aerial operating conditions.
Impact of Altitude and Mobility: Measurements collected at multiple altitudes and along controlled flight trajectories enable analysis of how UAV height and mobility influence propagation conditions and coverage. The dataset allows researchers to isolate altitude- and speed-dependent effects on channel behavior.
Waveform and System Design Evaluation: The raw I/Q samples and associated metadata support simulation and evaluation of waveform and receiver designs under realistic channel conditions. These data enable assessment of system performance using experimentally captured aerial signals rather than idealized channel assumptions.
Validation of Ray-Tracing and Analytical Models: The dataset provides empirical reference data that can be used to validate and calibrate ray-tracing simulations and analytical propagation models. Comparisons between measured and modeled results support assessment of model accuracy in rural air-to-ground environments.
@@ X Wireless Localization Dataset
(a) Keysight N6841A RF sensor deployed on AERPAW tower LW3.
(b) AERPAW UAV with SDR portable node.
(c) AERPAW’s LWRFL and RF sensor tower locations at LW2, LW3, LW4, LW5.
Figure 22: Overview of the AERPAW testbed showing the RF sensor, UAV node, and LWRFL site layout. (a) Keysight N6841A RF sensor deployed on AERPAW tower LW3, (b) AERPAW UAV with SDR portable node, and (c) AERPAW’s LWRFL and RF sensor tower locations at LW2, LW3, LW4, LW5.
@@ X-A Description of Hardware and Software
UAV flight experiments were conducted in AERPAW’s controlled environment to generate both datasets, using Keysight N6841A RF sensors to collect TDOA measurements. The N6841A (Fig. 22a) is capable of detecting, recording, and precisely time-stamping RF signals across a frequency range of 20 MHz to 6 GHz, with a maximum bandwidth of 20 MHz. Equipped with a broadband omnidirectional antenna and GPS-based timestamping, the system uses Keysight’s N6854A Geolocation Software and Sensor Management Tool to support TDOA, RSS, and hybrid localization methods for tracking RF sources within approximately a 2 km radius. However, the software is limited to 2D localization and does not estimate altitude.
Fig. 22c shows the deployment of these sensors at the LWRFL, where a single N6841A unit is mounted on each of the four towers labeled LW2 through LW5, approximately 10 meters above ground level. The sensors are deployed in a rural environment with mixed LoS conditions due to tree cover and building obstructions. All RF sensors within AERPAW are synchronized using a shared GPS-disciplined clock infrastructure and operate on the same local network, with individual IP addresses assigned within a common subnet. Centralized management via the Keysight Geolocation Server ensures network-level time synchronization, which is essential for accurate TDOA-based localization. During experiments, the N6841A units capture I/Q data from RF signals, which are subsequently processed to estimate UAV positions. However, the specific algorithms used for TDOA extraction and position estimation are proprietary to Keysight and not publicly available.
In our previous work [17], a 3.32 GHz channel sounding waveform was transmitted from a UAV-mounted SDR (Fig. 22c), and the RF sensors localized the UAV along multiple repeated trajectories. The dataset associated with this study is publicly available at [18]. In a separate study [bhattacherjee2022experimental], TDOA-based localization was performed using downlink control signals from a DJI Inspire 2 UAV operating in the 2.400–2.483 GHz ISM band with a 20 MHz bandwidth. The dataset corresponding to this experiment is also publicly available at [13].
@@ X-B Dataset Format
The TDOA-based UAV localization datasets associated with [17, 18] provide time-synchronized localization estimates and performance metrics collected during multiple UAV flight experiments on the AERPAW platform. Each dataset corresponds to an individual flight conducted at a specified altitude and signal bandwidth, enabling controlled evaluation of TDOA localization accuracy under varying propagation conditions.
For each flight, the dataset records estimated UAV positions together with ground-truth coordinates, center frequency information, timestamps, and localization performance indicators. These indicators include correlation-based metrics and error statistics that quantify localization accuracy relative to known ground truth. In addition, the datasets include binary indicators describing LoS or NLoS conditions between the UAV transmitter and multiple AERPAW RF sensor towers, enabling analysis of the impact of propagation conditions on localization performance.
A related set of TDOA localization datasets reported in [bhattacherjee2022experimental, 13] follows a similar structure but differs in the representation of LoS information and the separation of measurement and ground-truth data. Together, these datasets support benchmarking of TDOA localization algorithms, evaluation of LoS/NLoS effects, and comparative studies across experimental configurations.
The datasets are released in structured CSV format, along with auxiliary processing tools to support visualization and performance evaluation. Detailed file inventories, CSV field descriptions, and helper script information are provided in Appendix A.
@@ X-C Representative Results
(a) UAV trajectory at 40 meter altitude color-coded by LoS conditions to towers LW2-4.
(b) Ground truth UAV trajectory color-coded by localization error.
@@
Fig. 23a illustrates the UAV’s 40-meter altitude trajectory overlaid on the LWRFL site map, showing LoS conditions relative to three RF sensor towers (LW2, LW3, and LW4), which are indicated by green triangles. The trajectory is color-coded based on five distinct LoS scenarios, where each case indicates LoS or NLoS status relative to the three towers. For example, Case 1 (green) represents segments where the UAV maintained LoS to all three towers, while Case 5 (purple) corresponds to areas where only LW3 maintained LoS. The color-coding of the trajectory clearly highlights frequent transitions between LoS conditions as the UAV moves through the field, illustrating the mixed LoS/NLoS environment characteristic of the LWRFL site. A key takeaway from this figure is that LoS availability varies substantially along the trajectory, even at a fixed altitude, leading to position-dependent measurement quality for TDOA-based localization.
Fig. 23b presents a color-coded visualization of localization error along the UAV’s trajectory. Each point represents a ground truth UAV position, with circles indicating the corresponding localization error relative to the nearest estimated coordinate. Green circles mark locations where the position error is less than 100 meters, while red circles identify points where the error exceeds 100 meters. Blue triangles indicate the locations of the RF sensor towers used for TDOA-based localization.
As shown in the figure, the majority of the trajectory is associated with low localization error (green), suggesting consistent and accurate TDOA-based position estimates along most of the UAV’s flight path. Higher error regions (red) are concentrated in specific trajectory segments that coincide with reduced LoS availability (due to foliage obstructions) or unfavorable sensor geometry. Together, these results highlight two key observations: (i) reliable localization is achievable across most of the trajectory when favorable LoS conditions are available, and (ii) localization performance degrades in predictable regions where LoS conditions or geometric dilution of precision (GDOP) are poor. This emphasizes the strong coupling between propagation conditions, sensor geometry, and localization accuracy in real-world environments.
@@ X-D Possible Uses of Dataset
Tracking Filter Development and Evaluation: The dataset supports the development and evaluation of tracking filters, such as Kalman and particle filters, by providing real-world UAV trajectory and localization measurement sequences. These data enable state estimation studies under mixed line-of-sight and non-line-of-sight conditions.
TDOA Localization Algorithm Benchmarking: The dataset enables performance evaluation and benchmarking of TDOA-based localization algorithms using real-world RF measurements. It supports direct comparison against industry-standard implementations, including Keysight-based TDOA localization, under mixed LoS/NLoS conditions.
Multi-Sensor Fusion: The dataset supports the development and evaluation of multi-sensor fusion algorithms by combining TDOA-based RF localization data with complementary sensing modalities. Such integration enables improved UAV tracking and state estimation in complex propagation environments.
Non-Line-of-Sight Modeling and Analysis: The dataset enables characterization and modeling of NLoS effects by providing real-world localization errors alongside explicit LoS/NLoS labels. These data support studies on NLoS detection, mitigation strategies, and bias-aware localization techniques.
CRLB Modeling and Validation: The dataset supports theoretical performance analysis through CRLB modeling and enables empirical validation by comparing theoretical bounds against measured TDOA-based localization errors under varying sensor geometries and environmental conditions.